OpenStudy (anonymous):

Which is an equation in point-slope form for the given point and slope? Point: (–8, 3); Slope: 6 Will give medal!!!!

OpenStudy (midhun.madhu1987):

y - y1 = m(x - x1)

OpenStudy (anonymous):

well point slope formula is \[y-y _{1}=m(x-x _{1})\]

OpenStudy (midhun.madhu1987):

where m = slope

OpenStudy (anonymous):

the point slop formula is y-y1=m(x-x1). Just plug in for the given variables. Remember that m is the slope.

OpenStudy (anonymous):

hey im just studying this

OpenStudy (anonymous):

Math is fun!

OpenStudy (anonymous):

Point-slope form is as follows: (y-y(1))=m(x-x(1)) Where y(1) is the y component of a point, x(1) is the x component of a point, and m is the slope.

OpenStudy (anonymous):

so easy

OpenStudy (anonymous):

Indeed.

OpenStudy (anonymous):

i shall have my medal now

OpenStudy (anonymous):

But people learn faster than other. Be considerate.

OpenStudy (anonymous):

SHE LEFT

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Don't stress it. I didn't get a medal :)